# New Synthetic Goldmine: Hand Joint Angle-Driven EMG Data Generation Framework for Micro-Gesture Recognition

### v2

@@ Introduction
Electromyography (EMG)-based gesture recognition has become a key modality for human-computer interaction (HCI), owing to its noninvasive sensing and ability to capture subtle gesture movement. It has shown great promise in applications such as neural prosthetics, AR/VR interfaces, and wearable assistive technologies. However, practical deployment remains constrained by several critical challenges: the scarcity of large-scale and diverse EMG datasets, limited generalization across users and tasks, and the difficulty in modeling temporal and semantic consistency of EMG patterns (Eddy, Scheme, and Bateman 2023).
@@
A major bottleneck lies in the scarcity and structural limitations of existing EMG datasets. Publicly available datasets such as Ninapro (Atzori et al. 2012), CapgMyo (Dai et al. 2021), CSL-HDEMG (Amma et al. 2015), and SEEDS (Matran-Fernandez et al. 2019) are predominantly collected under predefined and isolated gesture protocols, offering limited degrees of freedom (DoFs) and failing to support natural transitions or combinations of DoFs commonly required in real-world control scenarios. This lack of flexibility restricts the datasets’ utility for modeling intuitive, user-driven interactions. Although HIT-SimCo (Yang and Liu 2021) attempts to address this issue by recording randomized combinations of wrist DoFs, it is constrained to single-session recordings, thereby limiting its applicability to studies involving cross-day variability, learning adaptation, and long-term stability. The recently introduced emg2pose dataset (Salter et al. 2024) significantly advances the state of EMG data collection by offering large-scale, multi-user, and multi-session recordings with synchronized hand pose annotations. It comprises over 370 hours of data from 193 subjects, covering 29 gesture stages under varied sensor placements and session conditions. Compared to earlier datasets, emg2pose provides high-resolution joint-angle trajectories across the entire hand and introduces benchmark splits to evaluate generalization between users, between sessions and between tasks. An additional concern is the significant imbalance between dynamic and isometric contractions, which can lead to biased model performance and reduced generalizability across different contraction types.
@@
On the algorithmic side, effective EMG pattern recognition remains highly dependent on capturing the complex temporal and semantic relationships inherent in muscle signals. While earlier studies employed traditional machine learning with handcrafted features (Phinyomark and Scheme 2018; Oskoei and Hu 2007, 2008), such methods often suffer from suboptimal feature representations and poor scalability. Although recent deep learning approaches (Gallon et al. 2024; Xiong et al. 2021; Chen et al. 2020) have improved recognition performance through end-to-end learning, several critical issues persist. Despite recent advances, EMG-based gesture modeling continues to face three interrelated challenges. First, many existing models struggle to capture the intricate temporal dependencies between joint kinematics and EMG activation, resulting in decoded or generated signals that lack semantic coherence and physiological plausibility (Tacca et al. 2024; Sitole and Sup 2023; Zhang et al. 2022). Second, the scarcity and uneven distribution of gesture data hinder model robustness, while commonly adopted augmentation strategies—such as random noise injection (Fatayer, Gao, and Fu 2022) or label-conditioned generation (Xiong et al. 2024)—lack the precision and coverage required to synthesize high-quality samples in underrepresented regions of the EMG feature space (Nasrallah et al. 2023; Jiang et al. 2021). While a few domain-specific methods have been proposed for ECG and EEG (Wang et al. 2025a; Ding et al. 2024; Wang et al. 2025b), they still lack explicit mechanisms tailored to EMG-specific challenges. Third, current discriminator-based evaluation frameworks rely on narrow, signal-level criteria and are ill-equipped to assess the fidelity of generated EMG signals across multiple dimensions, including amplitude continuity, temporal alignment, and waveform morphology (Farago et al. 2022; Zhao et al. 2024; Tsinganos et al. 2020).
@@
To this end, we propose SeqEMG-GAN, a sequence-driven conditional generative model that synthesizes high-fidelity EMG signals based on hand joint kinematic trajectories. Our key motivation is to leverage joint angle sequences—an interpretable and compact representation of movement—as a prior control signal to guide the generation of temporally aligned, physiologically consistent EMG sequences. The proposed framework comprises four core components: (1) a joint angle encoder to extract movement embeddings, (2) a latent context encoder incorporating a novel Ang2Gist unit to model fine-grained temporal dependencies between motion and muscle activation, (3) a deep convolutional generator to produce raw EMG waveforms, and (4) a multi-perspective discriminator that evaluates signal fidelity across amplitude, timing, and morphological perspectives. This design enables both targeted synthesis in data-scarce regions and robust generalization across gestures and users.
@@
To augment scarce EMG data, early attempts such as temporal warping (Gu et al. 2018) and noise injection (Stachaczyk, Atashzar, and Farina 2020) have been explored, but they lack semantic consistency. Transfer learning (Ameri et al. 2019; Wu, Yang, and Sawan 2023) alleviates inter-user variability, yet it relies on large-scale source data. Earlier biomechanical models (Kiguchi et al. 2008) map EMG to joint motion via physical equations, but their complexity and reliance on expert knowledge limit scalability. Statistical models like GMM and HMM (Yang, Pan, and Li 2017) impose strong assumptions, reducing generative diversity and generalization across poses and users.
@@
GAN-Based Methods: (Mendez, Lhoste, and Micera 2022) applied GANs to generate synthetic EMG signals for hand grasp classification, leading to improved classifier performance in low-resource settings. (Chen et al. 2022) employed DCGANs to augment multi-channel EMG data represented as grayscale images, achieving marginal but consistent gains in recognition accuracy. (Lin et al. 2023) proposed RoHDE, a GAN-based framework that introduces perturbation patterns into HD-EMG to enhance classifier robustness against real-world signal artifacts. (Lin et al. 2023) introduced a style-transfer GAN to disentangle identity-specific signal features from gesture-related information, highlighting privacy vulnerabilities and EMG anonymization opportunities.
Privacy-Preserving Synthesis: (Hazra and Byun 2020) designed SynSigGAN to generate synthetic biomedical signals, including EMG, achieving high signal correlation and enabling data sharing under privacy constraints.
Diffusion-Based Methods: Recent diffusion-based approaches have shown superior modeling of complex EMG distributions. (Xiong et al. 2024) introduced PatchEMG, which employs patch-wise diffusion for few-shot gesture recognition, demonstrating high fidelity synthesis with limited data. (Liu et al. 2024) proposed SDEMG, a score-based diffusion model that effectively removes ECG artifacts from diaphragmatic EMG. (Qing et al. 2024) utilized a diffusion model to synthesize realistic EMG for ergonomic studies, preserving utility while mitigating privacy concerns.
VAE-Based Methods:  (Olsson et al. 2021) propose a VAE-based framework for open-set gesture recognition using sparse multichannel EMG signals. Their method enables reliable classification of known gestures while rejecting unknown ones, addressing a key need in real-world prosthetic control. Unlike closed-set models, the VAE captures uncertainty, allowing effective separation of seen and unseen gestures.
Cross-Modal and Multimodal Generative Models based on EMG
Cross-modal generative models that involve EMG as either source or target modality have opened new directions for silent speech communication and biosignal augmentation.
EMG-to-Speech Translation: (Scheck and Schultz 2023a) proposed EMGVox-GAN, a GAN-based vocoder that synthesizes intelligible speech directly from facial EMG inputs, offering promise for silent speech interfaces and assistive technologies.
Speech-to-EMG Synthesis: (Scheck and Schultz 2023b) introduced STE-GAN to generate EMG signals from acoustic speech, facilitating augmentation of EMG datasets by leveraging larger speech corpora. This enables effective training of EMG-based speech recognizers with minimal biosignal recordings.
@@
Given a joint angle sequence $S=\{s_{1},s_{2},\dots,s_{T}\}$, the model generates a corresponding EMG sequence $\hat{X}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T}\}$. The generation process is governed by a global latent context vector $\mathbf{h}_{0}$, which summarizes high-level motion semantics. To enable stochastic generation and improve generalization, we apply the reparameterization trick (Kingma and Welling 2013) to sample $\mathbf{h}_{0}$ from a learned posterior distribution:
@@
  • Lower Layer: Standard GRU units encode motion dynamics.
  • Upper Layer: A novel Ang2Gist unit refines contextual representation.
@@
Here, $\sigma_{\{z,r,h\}}$ are nonlinear activations; $\mathbf{W}_{\ast}$, $\mathbf{U}_{\ast}$, and $\mathbf{b}_{\ast}$ are trainable parameters; $\odot$ and $\ast$ denote element-wise product and filtering, respectively.
@@
The experiments in this chapter utilize the latest emg2pose dataset proposed by Meta (Salter et al. 2024). This dataset is currently the largest publicly available high-quality dataset for wrist-based electromyography (EMG) recordings with hand gesture annotations. The emg2pose dataset includes 16-channel EMG signals sampled at 2 kHz, along with precise gesture annotations captured by 26 cameras. It covers a total of 193 users, 370 hours of recording time, and 29 gesture categories, ensuring a high degree of gesture diversity.
@@
To assess the effectiveness of SeqEMG-GAN, we conducted two main experiments: similarity analysis and performance validation. Given that standard GAN evaluation metrics such as Inception Score (IS), Frechet Inception Distance (FID), and Kernel MMD (Xu et al. 2018) are primarily designed for image generation and high-dimensional feature space comparison, they are not directly applicable to EMG time-series data.
To address this, we adopt metrics more suitable for temporal biomedical signals, following prior work (Delaney, Brophy, and Ward 2019), and propose a multi-perspective evaluation scheme covering frequency-domain, time-domain, and signal morphology consistency:
@@
To measure structural similarity in the time domain, we use DTW to non-linearly align generated and real sequences. This technique is especially robust to temporal shifts and warping in EMG signals. We implement FastDTW (Salvador 2007) to reduce computational cost to $O(N)$.
EMG Envelope Cross-Correlation (CC)
We evaluate global shape consistency using cross-correlation between smoothed signal envelopes. The envelope is computed by applying a moving average filter to the absolute EMG signals, and normalized CC quantifies trend alignment and tremor consistency.
@@
The baseline GAN model performs the worst across all metrics, failing to support conditional generation or unknown gesture generation. In 2020, Rafael Anicet Zanini et al. (Anicet Zanini and Luna Colombini 2020) proposed an EMG synthesis method by combining DCGAN with Style Transfer. Their approach employs a CNN-based architecture to generate EMG signals and introduces four feature pipelines (raw EMG, FFT, envelope FFT, and wavelet decomposition) to enrich the feature space, making the generated signals more realistic. While this method supports conditional generation, it lacks the ability to generate unknown gestures.
The DCGAN model, proposed by Yaojia Qian et al. (Wu, Chen, and Meng 2020), applies matrix transformation normalization to convert EMG signals into grayscale images before training a network using image generation techniques. Although this approach leverages advances in image synthesis, it suffers from information loss. Like the method of Zanini et al., DCGAN supports conditional generation but is incapable of generating unknown gestures.
@@
Table 1: Similarity Analysis Between Generated and Real Data
[…]
GAN $\times$ $\times$ 103.427321 19.564078 0.719453
StyleTransfer $\surd$ $\times$ 98.532786 13.531477 0.781721
DCGAN $\surd$ $\times$ 93.439412 9.675622 0.791930
Ours $\surd$ $\surd$ 91.756312 8.764323 0.817492
@@
To validate the effectiveness of the data generated by SeqEMG-GAN, we use classification accuracy as the evaluation metric. In particular, we randomly select six identical gesture categories from either generated or real data and use them as the training set to train a classifier. The classifier’s performance is then evaluated to determine whether the data contributes to model training. In this study, we select two traditional machine learning classification algorithms: Support Vector Machine (SVM) and Random Forest (RF), along with two deep learning-based classifiers. The experimental results are presented in Table 2.
@@
  • RR: The classifier is trained using 80% real data and tested on 20% real data.
@@
Table 2 presents the classification accuracy results for performance validation analysis. The results indicate that when classifiers are trained solely on synthetic data, the average accuracy decreases by 2.06%, dropping from 57.77% to 55.71%. However, by combining real data with synthetic data for training, the average accuracy significantly increases to 60.53%, achieving an improvement of 2.76% compared to classifiers trained solely on real data. This finding suggests that integrating synthetic data with real data can effectively enhance classifier performance.
@@
Removing the Angle Encoder (a) leads to a substantial drop in envelope correlation and an increase in DTW, highlighting the importance of high-level joint representation. Disabling the GRU module (b) causes a pronounced degradation in DTW and correlation, suggesting temporal modeling is essential for sequential consistency. Excluding the Ang2Gist blocks (c) reduces semantic alignment, as reflected in lower envelope correlation and worse frequency-domain similarity. Finally, the non-GAN baseline (d) results in notably smoother but less realistic signals, indicated by the highest FFT MSE and lowest envelope correlation, despite having slightly better DTW due to less fluctuation.

### v3

@@ Introduction
Electromyography (EMG)-based gesture recognition has become a key modality for human-computer interaction (HCI), owing to its noninvasive sensing and ability to capture subtle gesture movement. It has shown great promise in applications such as neural prosthetics, AR/VR interfaces, and wearable assistive technologies. However, practical deployment remains constrained by several critical challenges: the scarcity of large-scale and diverse EMG datasets, limited generalization across users and tasks, and the difficulty in modeling temporal and semantic consistency of EMG patterns (eddy2023framework).
@@
A major bottleneck lies in the scarcity and structural limitations of existing EMG datasets. Publicly available datasets such as Ninapro (atzori2012building), CapgMyo (dai2021capg), CSL-HDEMG (amma2015advancing), and SEEDS (matran2019seeds) are predominantly collected under predefined and isolated gesture protocols, offering limited degrees of freedom (DoFs) and failing to support natural transitions or combinations of DoFs commonly required in real-world control scenarios. This lack of flexibility restricts the datasets’ utility for modeling intuitive, user-driven interactions. Although HIT-SimCo (yang2021emg) attempts to address this issue by recording randomized combinations of wrist DoFs, it is constrained to single-session recordings, thereby limiting its applicability to studies involving cross-day variability, learning adaptation, and long-term stability. The recently introduced emg2pose dataset (salter2024emg2pose) significantly advances the state of EMG data collection by offering large-scale, multi-user, and multi-session recordings with synchronized hand pose annotations. It comprises over 370 hours of data from 193 subjects, covering 29 gesture stages under varied sensor placements and session conditions. Compared to earlier datasets, emg2pose provides high-resolution joint-angle trajectories across the entire hand and introduces benchmark splits to evaluate generalization between users, between sessions and between tasks. An additional concern is the significant imbalance between dynamic and isometric contractions, which can lead to biased model performance and reduced generalizability across different contraction types.
@@
On the algorithmic side, effective EMG pattern recognition remains highly dependent on capturing the complex temporal and semantic relationships inherent in muscle signals. While earlier studies employed traditional machine learning with handcrafted features (phinyomark2018emg; oskoei2007myoelectric; oskoei2008support), such methods often suffer from suboptimal feature representations and poor scalability. Although recent deep learning approaches (gallon2024comparison; xiong2021deep; chen2020surface) have improved recognition performance through end-to-end learning, several critical issues persist. Despite recent advances, EMG-based gesture modeling continues to face three interrelated challenges. First, many existing models struggle to capture the intricate temporal dependencies between joint kinematics and EMG activation, resulting in decoded or generated signals that lack semantic coherence and physiological plausibility (tacca2024wearable; sitole2023continuous; zhang2022ankle). Second, the scarcity and uneven distribution of gesture data hinder model robustness, while commonly adopted augmentation strategies—such as random noise injection (fatayer2022semg) or label-conditioned generation (xiong2024patchemg)—lack the precision and coverage required to synthesize high-quality samples in underrepresented regions of the EMG feature space (nasrallah2023semg; jiang2021data). Third, current discriminator-based evaluation frameworks rely on narrow, signal-level criteria and are ill-equipped to assess the fidelity of generated EMG signals across multiple dimensions, including amplitude continuity, temporal alignment, and waveform morphology (farago2022review; zhao2024trgan; tsinganos2020data).
@@
Joint angle–driven conditioning is a promising direction: joint angles summarize underlying activations, and conditioning on them renders EMG synthesis more plausible under inverse-dynamics constraints on feasible muscle activity. Here we introduce SeqEMG-GAN, a sequence-driven conditional generative model that synthesizes high-fidelity EMG signals based on hand joint kinematic trajectories. Our key motivation is to leverage joint angle sequences—an interpretable and compact representation of movement—as a prior control signal to guide the generation of temporally aligned, physiologically consistent EMG sequences. The proposed framework comprises four core components: (1) a joint angle encoder to extract movement embeddings, (2) a latent context encoder incorporating a novel Ang2Gist unit to model fine-grained temporal dependencies between motion and muscle activation, (3) a deep convolutional generator to produce raw EMG waveforms, and (4) a multi-perspective discriminator that evaluates signal fidelity across amplitude, timing, and morphological perspectives. This design enables both targeted synthesis in data-scarce regions and robust generalization across gestures and users.
@@
To augment scarce EMG data, early attempts such as temporal warping (gu2018robust) and noise injection (stachaczyk2020adaptive) have been explored, but they lack semantic consistency. Transfer learning (ameri2019deep; wu2023transfer) alleviates inter-user variability, yet it relies on large-scale source data. Earlier biomechanical models (kiguchi2008development) map EMG to joint motion via physical equations, but their complexity and reliance on expert knowledge limit scalability. Statistical models like GMM and HMM (yang2017semg) impose strong assumptions, reducing generative diversity and generalization across poses and users.
@@
GAN-Based Methods: (mendez2022emg) applied GANs to generate synthetic EMG signals for hand grasp classification, leading to improved classifier performance in low-resource settings. (chen2022deep) employed DCGANs to augment multi-channel EMG data represented as grayscale images, achieving marginal but consistent gains in recognition accuracy. (lin2023toward) proposed RoHDE, a GAN-based framework that introduces perturbation patterns into HD-EMG to enhance classifier robustness against real-world signal artifacts. (lin2023toward) introduced a style-transfer GAN to disentangle identity-specific signal features from gesture-related information, highlighting privacy vulnerabilities and EMG anonymization opportunities.
Privacy-Preserving Synthesis: Beyond task-driven synthesis, (hazra2020synsiggan) designed SynSigGAN to generate synthetic biomedical signals, including EMG, achieving high signal correlation and enabling data sharing under privacy constraints.
Diffusion-Based Methods: Recent diffusion-based approaches have shown superior modeling of complex EMG distributions. (xiong2024patchemg) introduced PatchEMG, which employs patch-wise diffusion for few-shot gesture recognition, demonstrating high fidelity synthesis with limited data. (liu2024sdemg) proposed SDEMG, a score-based diffusion model that effectively removes ECG artifacts from diaphragmatic EMG. (qing2024enhancing) utilized a diffusion model to synthesize realistic EMG for ergonomic studies, preserving utility while mitigating privacy concerns.
VAE-Based Methods: Latent-variable models offer calibrated uncertainty that is valuable for open-set scenarios.  (olsson2021can) propose a VAE-based framework for open-set gesture recognition using sparse multichannel EMG signals. Their method enables reliable classification of known gestures while rejecting unknown ones, addressing a key need in real-world prosthetic control. Unlike closed-set models, the VAE captures uncertainty, allowing effective separation of seen and unseen gestures.
@@
Given a joint angle sequence $S=\{s_{1},s_{2},\dots,s_{T}\}$, the model generates a corresponding EMG sequence $\hat{X}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T}\}$. The generation process is governed by a global latent context vector $\mathbf{h}_{0}$, which summarizes high-level motion semantics. To enable stochastic generation and improve generalization, we apply the reparameterization trick (kingma2015variational) to sample $\mathbf{h}_{0}$ from a learned posterior distribution:
@@
  • Lower Layer: Standard Gated Recurrent Unit (GRU) units encode motion dynamics.
  • Upper Layer: A novel Angle‑to‑Gist (Ang2Gist) unit refines contextual representation.
@@
Here, $\sigma_{\{z,r,h\}}$ are nonlinear activations; $\mathbf{W}_{\ast}$, $\mathbf{U}_{\ast}$, and $\mathbf{b}_{\ast}$ are trainable parameters; $\odot$ and $\ast$ denote element-wise product and filtering, respectively. Filter(·) is a 1D depthwise‑separable convolution along time (kernel size = 5, stride = 1, padding = 2) applied to the angle embedding $i_{t}$; its parameters are learned jointly with the generator (Xavier‑uniform initialization, SGD with momentum), and each filter channel is shared across gesture dimensions to enforce temporal consistency at low cost.
@@
The experiments in this chapter utilize the latest emg2pose dataset proposed by Meta (salter2024emg2pose). This dataset is currently the largest publicly available high-quality dataset for wrist-based electromyography (EMG) recordings with hand gesture annotations. The emg2pose dataset includes 16-channel EMG signals sampled at 2 kHz, along with precise gesture annotations captured by 26 cameras. It covers a total of 193 users, 370 hours of recording time, and 29 gesture categories, ensuring a high degree of gesture diversity.
@@
To assess the effectiveness of SeqEMG-GAN, we conducted two main experiments: similarity analysis and performance validation. Given that standard GAN evaluation metrics such as Inception Score (IS), Frechet Inception Distance (FID), and Kernel MMD (xu2018diversity) are primarily designed for image generation and high-dimensional feature space comparison, they are not directly applicable to EMG time-series data.
To address this, we adopt metrics more suitable for temporal biomedical signals, following prior work (delaney2019synthesis), and propose a multi-perspective evaluation scheme covering frequency-domain, time-domain, and signal morphology consistency:
@@
To measure structural similarity in the time domain, we use DTW to non-linearly align generated and real sequences. DTW is a distance between two time series that allows non-linear time stretching/compression (lower is better). This technique is especially robust to temporal shifts and warping in EMG signals. We implement FastDTW (salvador2007toward) to reduce computational cost to $O(N)$.
EMG Envelope Cross-Correlation ($C$)
We evaluate global shape consistency using cross-correlation between smoothed signal envelopes. The envelope is computed by applying a moving average filter to the absolute EMG signals, and normalized $C$ quantifies trend alignment and tremor consistency.
@@
[…]
GAN $\times$ $\times$ 103.427321 19.564078 0.719453
StyleTransfer $\surd$ $\times$ 98.532786 13.531477 0.781721
DCGAN $\surd$ $\times$ 93.439412 9.675622 0.791930
Ours $\surd$ $\surd$ 91.756312 8.764323 0.817492
Table 1: Similarity Analysis Between Generated and Real Data. ‘Conditional Generation’ = generation controllable via joint-angle input; ‘Unknown Gesture Generation’ = synthesis for gestures absent from the generator’s training set.
@@
The baseline GAN model performs the worst across all metrics, failing to support conditional generation or unknown gesture generation. In 2020, Rafael Anicet Zanini et al. (anicet2020parkinson) proposed an EMG synthesis method by combining DCGAN with Style Transfer. Their approach employs a CNN-based architecture to generate EMG signals and introduces four feature pipelines (raw EMG, FFT, envelope FFT, and wavelet decomposition) to enrich the feature space, making the generated signals more realistic. While this method supports conditional generation, it lacks the ability to generate unknown gestures.
The DCGAN model, proposed by Yaojia Qian et al. (wu2020dcgan), applies matrix transformation normalization to convert EMG signals into grayscale images before training a network using image generation techniques. Although this approach leverages advances in image synthesis, it suffers from information loss. Like the method of Zanini et al., DCGAN supports conditional generation but is incapable of generating unknown gestures.
@@
To validate the effectiveness of the data generated by SeqEMG-GAN, we use classification accuracy as the evaluation metric. In particular, we select six gestures from either generated or real data and use them as the training set to train a classifier. We classify Slide Left/Right/Up/Down, Click, Double Click — a standard HCI micro-gesture set aligned with widely recognized EMG wristband interfaces, ensuring real-world applicability. The classifier’s performance is then evaluated to determine whether the data contributes to model training. In this study, we select two traditional machine learning classification algorithms: Support Vector Machine (SVM) and Random Forest (RF), along with two deep learning-based classifiers. The experimental results are presented in Table 2.
@@
  • $R$: The classifier is trained using 80% real data and tested on 20% real data.
@@
Table 2: Classification Accuracy Analysis
Table 2 presents the classification accuracy results for performance validation analysis. The results indicate that when classifiers are trained solely on synthetic data, the average accuracy decreases by 2.06%, dropping from 57.77% to 55.71%. However, by combining real data with synthetic data for training, the average accuracy significantly increases to 60.53%, achieving an improvement of 2.76% compared to classifiers trained solely on real data. Qualitatively, Slide gestures exhibit smoother envelope transitions than Click/Double-Click, aligning with their kinematic phases and supporting our angle-conditioned design. Taken together, these results indicate that integrating synthetic with real data effectively enhances classifier performance.
@@
Cross‑user/session Generalization
Surface EMG varies markedly across users (anatomy, muscle recruitment) and sessions (electrode shift, skin impedance, fatigue); thus, cross-subject/session evaluations directly test robustness to these distribution shifts.
Leveraging emg2pose metadata (held_out_user, held_out_stage, generalization), we enforce disjoint train/test partitions and evaluate four protocols: sample-wise, cross-subject, cross-session, and cross user+stage. Unless noted otherwise, models are trained under Real-only and Real+Synthetic (1:1) conditions, and results are reported as mean $\pm$ SD over five runs. Mixing synthetic data yields consistent gains across all protocols.
Setting Sample-wise Cross-Subj Cross-Sess Cross User+Stage
Real-only 57.8 $\pm$ 1.0 52.6 $\pm$ 1.1 56.0 $\pm$ 0.8 48.9 $\pm$ 1.2
Real + Synthetic 60.5 $\pm$ 0.9 57.1 $\pm$ 0.9 59.8 $\pm$ 0.7 54.3 $\pm$ 1.0
Table 3: Cross-user/session generalization accuracy (%). Adding synthetic data (1:1 mix). Means $\pm$ SD over five runs.
@@
Removing the Angle Encoder (a) leads to a substantial drop in envelope correlation and an increase in DTW, highlighting the importance of high-level joint representation. Disabling the GRU module (b) causes a pronounced degradation in DTW and correlation, suggesting temporal modeling is essential for sequential consistency. Excluding the Ang2Gist blocks (c) reduces semantic alignment, as reflected in lower envelope correlation and worse frequency-domain similarity. Finally, in the non-GAN baseline (d), DTW decreases slightly without the discriminator, but spectral and envelope consistency degrade markedly; overall fidelity is therefore highest for the full model.